A charge of -4 nC is located at (0,1.00) m. What is the x-component of the electric field at
(4.00, -2.00) m?
Answer #1
GivenMagnitude of the charge ,q = -4 nC at (0, 1.0m)The x component ofelectric field at (4.0 m , -2.0m) isEx = k q / r^2 cos36.86 [where r = √4^2 +3^2 = 5.0m]= (9.0 *10^9 Nm^2/C^2) * (4 *10^-9 C) * 0.8 /25 m^2= 1.15 N/C (-i)
answered by: La